# Geodetic Exploration of the Interior Structure of Enceladus

### v1

@@ 6.1 Constraining the interior and discriminating between heat generation scenarios
The tighter constraints obtained from Park et al. (2024) allow relaxation of the demands on measuring $h_{2}$ and $l_{2}$ in order to discriminate between the two heating scenarios. Based on the observational constraints of Park et al. (2024) (as in Figure 9a), we show the amplitudes and phase lags of the Love numbers in Figure 10 panels (a)-(f). Using this narrower set of observational constraints, the PDFs of the tidal response associated with the two tidal dissipation scenarios are further separated compared to Figure 8 where broader ranges of values for the observables were assumed. In this case, measuring the maximum vertical deformations with a precision of 87 cm is sufficient to differentiate between the two hypotheses of heat generation with a 1-$\sigma$ confidence. For a confidence level of 2-$\sigma$, measuring the maximum radial displacement of 1 cm is required. Alternatively, measuring the phase lag $\epsilon_{h_{2}}$ by 20 and 15 is sufficient to discriminate between the two tidal dissipation scenarios with 1-$\sigma$ and 2-$\sigma$ confidences, respectively. Also, our analysis indicates that measurements of maximum tangential displacements within 40 cm and 0.5 cm are required to differentiate between the two tidal dissipation scenarios providing 1-$\sigma$ and 2-$\sigma$ level confidences, respectively. Similarly to $\epsilon_{h_{2}}$, the measurement precisions required for $\epsilon_{l_{2}}$ for the confidence levels of 1-$\sigma$ and 2-$\sigma$ are 20 and 15, respectively.
@@
In case future measurements provide constraints on $k_{2}$, the measurement requirement for $h_{2}$ and/or $l_{2}$ could be significantly relaxed. Therefore, we present PDFs of displacement tidal Love numbers associated with the core- and shell-dominated tidal dissipation assuming additional constraints on the amplitude and phase lag of $k_{2}$ (Figure 10 panels (g)-(j)). If the real and imaginary parts of $k_{2}$ are measured within 0.002, the accuracies required to measure the maximum radial deformations to obtain 1-$\sigma$ and 2-$\sigma$ confidence levels relax to 134 cm and 82 cm, respectively. In that case, the upper bounds of the uncertainty required for the horizontal deformations are 61 cm and 36 cm, respectively. The accuracies required for the phase lags of $h_{2}$ or $l_{2}$ are both equal to 21 and 16, for confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively.
@@
Furthermore, in case a more precise measurement of $k_{2}$ (both amplitude and phase) is obtained, i.e. within an uncertainty of 0.0002, the required measurement of radial or horizontal deformations will be further relaxed (not shown here). In that case, the requirement for the maximum radial displacement will be 150 cm and 116 cm for the 1-$\sigma$ and 2-$\sigma$ confidence levels, respectively. These values are 68 cm and 51 cm for the tangential component of the deformation, respectively. The precisions required for the phase lags of $h_{2}$ or $l_{2}$ are equal to 24 and 22 to obtain confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively. We summarize the precision required for measuring the radial and horizontal displacements based on different scenarios discussed here in Table 5. Note that the measurements of the $h_{2}$ and $l_{2}$ as well as the $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ are correlated with each other, and therefore, measuring each of these four quantities are sufficient for the purpose of discriminating between tidal dissipation scenarios. We perform our analysis based on both the derivation in this paper (Equations 18) and that in Beuthe (2019). We find that the difference in the results is less than 5%. This close approximation indicates the applicability of (Equations 18) in capturing the tidal dissipation in the shell without a direct requirement of a priori constraint on the shell’s rheology.
Current Knowledge $k_{2}$ $\max\Delta R$ (cm) $\max\Delta E$ (cm) $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ (deg)
@@
Table 5: Range of the required precision for measuring $h_{2}$, $l_{2}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$ for different scenarios of available and future measurements. “Current knowledge” implies the use of $\gamma_{I}$, $J_{3}$, and heat loss from the referenced studies. “All studies” refers to the use of estimates covering Howett et al. (2011), Thomas et al. (2016), Nadezhdina et al. (2016) , Hemingway & Mittal (2019), and Park et al. (2024). In the two bottom cases, measurement uncertainties on $k_{2}$ are assumed and other requirements are computed correspondingly. The requirements are given for obtaining 2-$\sigma$ confidence on the dominating region of the heat dissipation (with those associated with the 1-$\sigma$ confidence provided in parenthesis). N/A indicates that given the corresponding uncertainties, obtaining the 2-$\sigma$ confidence is not possible.
@@
Figure 9a: Exploration of observables constraints from $J_{3}$, moment of inertia, libration amplitude, and heat loss rate based on Park et al. (2024). We presented PDFs of predictions for $k_{2}$, $h_{2}$, $l_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$. Both scenarios of tidal dissipation are considered here. The plots of the PDFs for the $k_{2}$ amplitude and phase lag are notably different than those of the $h_{2}$ and $l_{2}$ amplitudes and phases. Unit conventions are the same as in Figure 4.
Figure 9b: Exploration of the correlation between the model parameters associated with Figure 9a. Shear moduli are presented in GPa and other parameters are presented in their SI units. The ranges of the plausible interior parameters are shown in Table 4. Ice shell thickness, and shear moduli is plotted in logarithmic axes.
@@
Figure 10: Panels (a)-(f): Tidal Love numbers and phase lags for heat dissipation in the shell and in the core based on the constraints by Park et al. (2024). Panels (g)-(j): Love numbers and phase lags for heat dissipation in the shell and in the core assuming that the real and imaginary parts of $k_{2}$ are measured within 0.002.
@@ 6.2 Constraining the interior structure with future observations
We use the favored values of the observables obtained in Section 6.1 along with the uncertainty of the measurements expected from future missions to constrain the main interior structure parameters. Due to the previously highlighted non-uniqueness of the solutions relying solely on the measurement of $k_{2}$, we perform two separate exemplar inversions associated with heat generation in the ice shell or in the core. The central value of the observation and the uncertainties adopted in the inversion are shown in Table 3. We use the favored values given in this table for $J_{3}$, contrary to the mean value of 181.066$\times$10-6 (adjusted for a reference radius of 251.99 km used here as compared to Park et al. (2024)). Our preferred value $J_{3}$ is within the 2-$\sigma$ uncertainty range of that suggested by Park et al. (2024). We use the lowest necessary precision for measuring $h_{2}$ to discriminate between the two scenarios of heat dissipation derived in Section 6.1. We conduct the inversion using MCMC sampling and the synthetic observations for two scenarios of measurement uncertainty given in Table 3 and discussed in Section 6.4. Each of our MCMC simulations includes approximately 600,000 accepted models which ensures that the parameter space is appropriately explored (algorithm details in Appendix C). The posterior distributions of the model parameters are demonstrated in Figures 11 and 12 for the two sets of inversions (only shown for the higher uncertainty case). Results based on both cases of higher and lower uncertainties are summarize in Table 6.
For the case of lower measurement precisions in Table 3, the results show that the suite of observations can constrain the thickness of the shell and the radius of the rocky core to within $\pm$2.7 km, and $\pm$7 km, respectively (3-$\sigma$ confidence level). When tighter constraints on the observables are adopted, the mean thickness of the shell can be constrained within approximately 1.2 km. The ocean density is constrained within $\pm$60 kg/m3 and $\pm$45 kg/m3 in the cases where the tidal dissipation is dominated in the core and the shell, respectively. This low precision is roughly equivalent to a knowledge of seawater salinity to within 40 psu (Millero et al., 2008, Earth’s ocean salinity is 35 psu or parts per thousand). An improved resolution of approximately $\pm$20 kg/m3 can be obtained if the higher precision scenario mentioned in Table 3 is used. The corresponding 20 psu salinity precision would allow for distinguishing the nominal 10 psu salinity of Postberg et al. (2009b) from an earthlike or greater salinity.
@@
In the case of a tidally active shell, the viscosity of the ice shell is constrained in the range of 4.4$\times$1013–9.4$\times$1013 Pa.s to satisfy the tidal heating constraint. In this tidal dissipation scenario, the shear modulus of the shell is constrained by the tidal Love numbers to be within 2.3-3.4 GPa. When tidal dissipation is dominated in the core, the shell rheology remains largely unconstrained, due to the insensitivity of the tidal heating to ice rheology in this scenario. We constrain the effective viscosity of the core in this scenario to be 9.1$\times$1011–1.4$\times$1012 Pa.s. A tighter constraint on the core viscosity is obtained (9.5$\times$1011–1.2$\times$1012 Pa.s), if the higher precision on the measurements are employed.In contrast to the viscosity of the core, its rigidity remains largely unconstrained.
Figure 11: Results of the MCMC inversion for the case where the tidal dissipation is dominantly in the core, provided the uncertainty on the measurements as shown in Table 3. The shear moduli are plotted in log axes. Results are presented for two sets of uncertainties presented Table 3.
Figure 12: Results of the MCMC inversion for the case where the tidal dissipation is dominantly in the ice shell and adopting the measurement uncertainty given in Table 3. The shear moduli are plotted in log axes. Results are presented for two sets of uncertainties presented Table 3.
@@
$\eta_{\rm core}$ Pa.s 9.1$\times$ 1011–1.4$\times$1012 (9.5$\times$1011–1.2$\times$1012) 1016–1020 (1016–1020)
Table 6: Posterior ranges of the model parameters from two exemplars of MCMC inversions using synthetic future measurements with expected accuracies presented in Table 3. Ranges are presented for two scenarios of measurement uncertainties.
6.3 Structural Heterogeneity in the ice shell
We assess the effect of lateral variations in ice shell thickness on the tidal response and its effect on the required measurements of $k_{2m}$ and $h_{2m}$. We build on the analysis of Berne et al. (2023) and conduct simulations of tidal deformation for four classes of models: 1. without lateral variations in ice shell thickness; 2. with lateral variations in ice shell thickness; 3. with a dissipative core and no lateral variations in ice shell thickness; and 4. with a dissipative core and lateral variations in ice shell thickness. For each model class, we vary the mean ice shell thickness across 15–50 km consistent with the finding based on the available measurements discussed in Section 6.1. We assume a core of viscosity $10^{20}$ Pa.s for the case of a non-dissipative core, and $10^{12}$ Pa.s for the case of a dissipative core. All other parameters are chosen to be the same as those used in Figures 1 and 2. To compute $k_{2m}$ and $h_{2m}$, we use a numerical finite element model (Berne et al., 2023). To generate crustal thickness variations, we apply the topography model in Park et al. (2024) (with minor modifications) evaluated up to $l=8$ to the outer surface of geometries and compensate non-hydrostatic components of this structure at the ice-ocean boundary based on the formulation for Airy-type isostasy described in Hemingway & Matsuyama (2017). The shape model suggests significant variations in the thickness of the shell (Figure 14), ranging between 5 and 50 km.
@@ 6.3 Structural Heterogeneity in the ice shell
Figure 13: Top: gravitational Love numbers, $k_{20}$ (left column) and $k_{22}$ (right column), against mean ice shell thickness for models that assume a spherically symmetric ice shell (dotted line), an ice shell with lateral variations in thickness (solid line), or a dissipative core (dotted and solid red lines). Bottom: Similar to top row except showing displacement Love numbers $h_{20}$ and $h_{22}$ plotted against mean ice shell thickness. x-axes are plotted on a log-scale.
We also assess the effect of the ice shell’s structural heterogeneity on the maximum radial and horizontal tidal deformations on the surface compared to those derived in Section 6.1. The overall maximum surface deformation can differ in comparison to the spherically symmetric case used in previous studies. We compute the maximum displacements on the radial, northerly, and easterly directions for an ice shell with variable thicknesses and a spherically symmetric ice shell of the same average thickness equal to 30 km (Figure 15). In Figure 15, it is assumed that the tidal dissipation occurs in the core, and the viscosity of the ice shell is sufficiently high to assume a purely elastic behavior. The ice shell with heterogeneous thickness demonstrates variations in the radial and horizontal directions reaching 50 cm compared to a spherically symmetric shell (Figure 15). Larger deviations are observed for all components around the South pole, as expected, due to the significant variations in the thickness of the shell. The overall maximum radial displacement for the case of 3D structure is approximately 20 cm higher than the spherically symmetric case. The significant deviation of surface deformation demonstrated in these plots implies that for Enceladus, $h_{2}$ and $l_{2}$ can provide rough estimates of the actual surface displacements as commonly used for spherically symmetric bodies, and for a precise calculation of the displacement at any point on the surface the effect of the 3D structure is important. Since the globally maximum radial and horizontal displacements for a spherically symmetric body are smaller than those for a shell with 3D structural variations, the required accuracies presented in Section 6.1 derived by assuming spherical symmetry are conservative and thus remain valid.
@@
Figure 14: Variation of Enceladus’s ice shell thickness based on the shape model by Park et al. (2024) assuming isostasy. Contours denote intervals of 2 km.
@@
Figure 15: Amplitude of tidal deformation in radial, northerly, and easterly directions for two cases of an ice shell with variable thickness as shown in Figure 14 and a spherically symmetric shell with the same average thickness of 30 km. Here, it is assumed that the tidal dissipation is happening in the rocky core and the viscosity of ice shell is 1020 Pa.s, implying elastic deformations in the shell. Contours denote intervals of 0.05 m

### v3

@@ 6.1.1 Plausible Interior Parameters and Observables
The tighter constraints obtained from Park et al. (2024) allow relaxation of the measurement requirement for $h_{2}$ and $l_{2}$ in order to discriminate between the two heating scenarios. Based on the observational constraints of Park et al. (2024), we show the amplitudes and phase lags of the Love numbers in Figure 8 panels (a)–(f). Using this narrower set of observational constraints, the PDFs of the tidal response associated with the two tidal dissipation scenarios are further separated compared to Section 6.1.1, where broader ranges of values for the observables were assumed. In this case, measuring the maximum vertical deformations with a precision of 87 cm is sufficient to differentiate between the two hypotheses of heat generation with a 1-$\sigma$ confidence. For a confidence level of 2-$\sigma$, measuring the maximum radial displacement of 1 cm is required. Alternatively, measuring the phase lag $\epsilon_{h_{2}}$ by 20 and 15 is sufficient to discriminate between the two tidal dissipation scenarios with 1-$\sigma$ and 2-$\sigma$ confidences, respectively. In addition, our analysis indicates that measurements of the maximum tangential displacements within 40 cm and 0.5 cm are required to differentiate between the two tidal dissipation scenarios providing 1-$\sigma$ and 2-$\sigma$ level confidences, respectively. Similarly to $\epsilon_{h_{2}}$, the measurement precisions required for $\epsilon_{l_{2}}$ for the confidence levels of 1-$\sigma$ and 2-$\sigma$ are 20 and 15, respectively.
@@
In case future measurements provide constraints on $k_{2}$, the measurement requirement for $h_{2}$ and/or $l_{2}$ could be significantly relaxed. We present PDFs of displacement tidal Love numbers associated with the core- and shell-dominated tidal dissipation assuming constraints on the amplitude and phase lag of $k_{2}$ (Figure 8 panels (g)–(j)). If the real and imaginary parts of $k_{2}$ are measured within 0.002, the accuracies required to measure the maximum radial deformations to obtain 1-$\sigma$ and 2-$\sigma$ confidence levels relax to 134 cm and 82 cm, respectively. In that case, the upper bounds of the uncertainty required for the horizontal deformations are 61 cm and 36 cm, respectively. The accuracies required for the phase lags of $h_{2}$ or $l_{2}$ are equal to 21 and 16 for confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively.
@@
Furthermore, in case more precise measurements of $k_{2}$ (both amplitude and phase) are obtained, i.e. within an uncertainty of 0.0002, the required measurement of radial or horizontal deformations will be further relaxed (not shown here). In that case, the requirement for the maximum radial displacement will be 150 cm and 116 cm for the 1-$\sigma$ and 2-$\sigma$ confidence levels, respectively. These values are 68 cm and 51 cm for the tangential component of the deformation, respectively. The precisions required for the phase lags of $h_{2}$ or $l_{2}$ are equal to 24 and $2$ to obtain confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively. We summarize the precision required for measuring the radial and horizontal displacements based on different scenarios discussed here in Table 5. Note that the measurements of the $h_{2}$ and $l_{2}$ as well as the $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ are correlated with each other and, therefore, measuring each of these four quantities is sufficient for discriminating between tidal dissipation scenarios. We perform our analysis based on both the derivation in this paper (Equations 16) and that in Beuthe (2019). We find that the difference in the results is less than 5%. This close approximation indicates the applicability of (Equations 16) in capturing the tidal dissipation in the shell without a direct requirement of a priori constraint on the shell’s rheology.
Current Knowledge $|k_{2}|$ $\max\Delta R$ (cm) $\max\Delta E$ (cm) $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ (deg)
@@
Table 5: Range of the required precision for measuring $h_{2}$, $l_{2}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$ for different scenarios of available and future measurements. “Current knowledge” implies the use of $\gamma$, $J_{3}$, and heat loss from the referenced studies. “All studies” refers to the use of estimates covering Howett et al. (2011), Thomas et al. (2016), Nadezhdina et al. (2016) , Hemingway & Mittal (2019), and Park et al. (2024). In the two bottom cases, measurement uncertainties on $k_{2}$ are assumed and other requirements are computed correspondingly. The requirements are given for obtaining 2-$\sigma$ confidence on the dominating region of the heat dissipation (with those associated with the 1-$\sigma$ confidence provided in parenthesis). N/A indicates that given the corresponding uncertainties, obtaining the 2-$\sigma$ confidence is not possible.
@@
Figure 7: Exploration of observables constraints from $J_{3}$, moment of inertia, libration amplitude, and rate of heat loss based on Park et al. (2024). We presented PDFs of predictions for $|k_{2}|$, $|h_{2}|$, $|l_{2}|$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$. Both scenarios of tidal dissipation are considered here. The plots of the PDFs for the $k_{2}$ amplitude and phase lag are notably different than those of the $|h_{2}|$ and $|l_{2}|$ amplitudes and phases. Plotting and unit conventions are the same as in Figure 4.
@@
Figure 8: Panels (a)–(f): Tidal Love numbers and phase lags for heat dissipation in the shell and in the core based on the constraints by Park et al. (2024). Panels (g)–(j): Love numbers and phase lags for two dissipation regimes assuming that the real and imaginary parts of $k_{2}$ are measured with an uncertainty of 0.002.
Figure 9a: Sensitivity of key parameters in determining the heat dissipation consistent with the observables in Figure 7 based on the analysis of the observables by Park et al. (2024). Rate of heat loss and $\mu_{shell}$ are shown in in GW and GPa, respectively. Blue and red colors demonstrate lowest and highest likelihoods, respectively.
[…]
Using the ranges of $\gamma$, $J_{3}$, MoI, and heat inferred by Park et al. (2024), we also obtain considerably improved constraints on the interior parameters compared to the case where we use broader uncertainties on the observations (See Figures 18 and 18a in Appendix E). However, several model parameters remain weakly constrained by the currently available observations. The ranges of the obtained plausible interior parameters are listed in Table 4. See Appendix E (Figures 18b and 18c) for the PDFs of all the model parameters associated with each of the cases. We further investigate the correlation between total heat and the observables and the interior parameters to which they are sensitive (Figure 9a). The ranges of the effective viscosities of the shell and the core are demonstrated. The total tidal heating is largely determined by the viscosity of the region where the tidal dissipation is dominated. The rigidity of the shell influences the amplitude of both $k_{2}$ and $h_{2}$. We also explore the correlation between the physical properties of the shell (thickness, density, and rigidity) and the observables that are most sensitive to them, i.e., $\gamma$, $J_{3}$, and $k_{2}$ (Figure 9b). Shell thickness and density influence both $\gamma$ and $J_{3}$, although the effect of the density on $J_{3}$ is stronger than on $\gamma$. $D_{shell}$ also strongly correlates with $k_{2}$. The effect of the shell rigidity on the libration amplitude is small compared to the thickness and density which determine the shell’s moment of inertia. This sensitivity suggests that the shell’s response to the libration forces is mainly determined by inertial forces, consistent with Hemingway & Nimmo (2024). The analysis shows that although libration is currently an effective constraint on the shell thickness, for obtaining very high accuracy on the shell and ocean thickness, measurements of the gravity field alongside the shape are much more impactful. Lower sensitivity of libration is largely due to its dependence on the moment of inertia and flexibility of the shell which include effects from thickness, density, and rigidity of the shell.
@@ 6.2 Constraining the Interior Structure with Future Observations
[…]
For the case of lower measurement precisions in Table 3, the results show that the suite of observations can constrain the thickness of the shell and the radius of the rocky core to within $\pm$2.7 km, and $\pm$7 km, respectively (3-$\sigma$ confidence level). When tighter constraints on the observables are adopted, the mean thickness of the shell can be constrained within approximately 1.5 km. The ocean density is constrained within $\pm$60 kg/m3 and $\pm$45 kg/m3 in the cases where the tidal dissipation is dominated in the core and the shell, respectively. This low precision is roughly equivalent to a knowledge of seawater salinity to within 40 psu (Millero et al., 2008, Earth’s ocean salinity is 35 psu or parts per thousand). An improved resolution of approximately $\pm$20 kg/m3 can be obtained if the higher precision scenario mentioned in Table 3 is used. The corresponding 20 psu salinity precision would allow for distinguishing the nominal 10 psu salinity of Postberg et al. (2009) from an Earth-like or greater salinity.
@@
[…]
Figure 10a: Results of the $MC$ inversion for the case where the tidal dissipation is dominantly in the ice shell and adopting the measurement uncertainty given in Table 3. The shear moduli are plotted in log axes. Results are presented assuming the larger set of uncertainties shown in Table 3.
Figure 10b: Results of the $MC$ inversion for the case where the tidal dissipation is dominantly in the core, provided the uncertainty on the measurements as shown in Table 3. The shear moduli are plotted in log axes. Results are presented assuming the larger set of uncertainties shown in Table 3.
@@
Table 6: Posterior ranges of the model parameters from two exemplars of $MC$ inversions using synthetic future measurements with expected accuracies presented in Table 3. Ranges are presented for two scenarios of measurement uncertainties.
6.3 Effect of Structural Heterogeneity in the Ice Shell
The required measurement precisions computed in Section 6.1 are based on a body with variable properties only in the radial direction (Equations 2.1). Based on shape and gravity observations, Enceladus’s ice shell deviates from spherical symmetry, which in turn can change its tidal response (Běhounková et al., 2017; Berne et al., 2023a, b). Here, we compute the tidal displacements of a heterogeneous ice shell and compare it to the response from a shell with constant thickness with same mean thicknesses. We compute the crustal thickness variations by applying the topography model in Park et al. (2024) (with minor modifications) evaluated up to $l=8$ to the outer surface of geometries and altering the shape of the ice-ocean boundary to mirror the non-hydrostatic components of topography at the outer surface. The relationship between structure at the ice-ocean boundary and surface topography follows a formulation for Airy-type isostasy (Hemingway & Matsuyama, 2017) (see e.g. Equation 19 in Berne et al. (2024)). The shape model suggests variations in the thickness of the shell ranging between 5 and 50 km (Figure 11).
[…]
@@ 6.3 Effect of Structural Heterogeneity in the Ice Shell
Figure 11: Variation of Enceladus’s ice shell thickness based on a modified shape model from Park et al. (2024) assuming isostasy. Contours denote intervals of 2 km.
We assess the effect of the ice shell’s structural heterogeneity on the maximum radial and horizontal tidal deformations on the surface compared to those derived in Section 6.1. The overall maximum surface deformation can differ in comparison to the case of constant thickness shell used in previous studies. We compute the maximum displacements in the radial, northerly, and easterly directions for an ice shell with variable thickness and an ice shell of constant thickness with the same average thicknesses equal to 30 km (Figure 12). In this figure, it is assumed that tidal dissipation occurs in the core and the viscosity of the ice shell is high enough to assume purely elastic behavior. The ice shell with heterogeneous thickness demonstrates variations in the radial and horizontal directions reaching 50 cm compared to a shell of constant thickness (Figure 12). Larger deviations are observed for all components around the South pole, as expected, due to the significant variations in the thickness of the shell (Berne et al., 2023b). The global maximum radial displacement for the case of a shell with variable thickness is approximately 20 cm higher than the constant thickness case. The significant deviation of surface deformation demonstrated in these plots implies that for Enceladus, $h_{2}$ and $l_{2}$ can provide rough estimates of the actual surface displacements as commonly used for spherically symmetric bodies, and for a precise calculation of the displacement at any point on the surface, the effect of structural heterogeneity is important. Since the globally maximum radial and horizontal displacements for a body of constant shell thickness are smaller than those for a shell with thickness variations, the required accuracies presented in Section 6.1 derived by assuming variation only in radial direction are conservative and thus remain valid.
@@
Figure 12: Amplitude of tidal deformation in radial, northerly, and easterly directions for two cases of an ice shell with variable thickness as shown in Figure 11 and a constant shell thickness with the same average thickness of 30 km. Here, it is assumed that the tidal dissipation is happening in the rocky core and the viscosity of ice shell is 1020 Pa.s, implying elastic deformations in the shell. Contours denote intervals of 0.05 m